For the following cashflows, the project may have up to?
Year Cashflow
0 -$500
1 0
2 0
3 400
4 -50
5 300

a. one RORs
b. two RORs
c. three RORs
d. Four RORs
1
Expert's answer
2015-11-03T00:00:47-0500
The rate of return is always expressed as a percentage and is calculated using the following formula:
ROR = (Net Income / Investment Cost) x 100
In our case maximal net income will be the sum of all cash flows with the 0% discount rate, so the maximal ROR is:
RORmax = (0 + 0 + 400 - 50 + 300)/500 = 650/500 = 1.3, so the right answer is a. one RORs.